OpenStudy (anonymous):

what steps would I use to solve the equation 1.8=(1.05)(power of X)

OpenStudy (anonymous):

Take the ln of both sides first and you will be able to apply log rules: \[ln(1.8)=ln(1.05^x)\] Per the log rule: \[ln(m^n)=nln(m)\] \[ln(1.8)=xln(1.05)\] \[x=\frac{ln1.8}{ln1.05}\] \[x=12.05\]

OpenStudy (anonymous):

Logarithms!!! log(1.8) = xlog(1.5) divide both side by log(1.5) x = 12.0472

OpenStudy (anonymous):

Alternatively you can also use log instead of ln.

OpenStudy (anonymous):

log = log base 10 ln = log base e use the change of base formulas to convert. log base x of y = log(y)/log(x)